Show full handtracing of the following code and describe what it accomplishes.
The input file, dataofUnknownorigin.in, contains the following data 11-8 2 0 0 6 3 include kiostream include <fstream using namespace std; struct a Thing double one: double two; bool theseAreGreenEggsAndHama aThing, aThing); (aThing, Thing from Horton (ifstreams); void drummingonADrum (aThing) int main if stream somewho; some Who.open (dataOfUnknownOrigin.in) a Thing thingon thing Two, thingThr thing one from Horton (somewho) while some ho.. eof thing Two fromHorton (somewho) thing Three fromHorton (somewho) if (the seareGreenEggsAnd am (thingone, thingTwo, thingThree)) cout YES else cout NO drummingonADrum (thingone); cout drumming OnADrum (thing Two); cout drummingonADrum (thingThree); cout endl thing one fromHorton (somewho); return 0; bool these AreGreenEggsAndHam (aThing tha, aThing thB, aThing thc) return (((thA. one-thC.cne)/ (thA two-thC-theo)) (thA. one-thB. one) (thA.two-thB., two))); a Thing fromHorton (ifstreams cindyLouiho) a Thing thingIt cindy Louw ho thingIt.one thing It.two; return thingIt; void drummingOnADrum CaThing oneThumb) cout oneThumb.one one Thumb. two
Answer #1

each line of file contains 6 inpus, (3 points in 2D plane).
   1) While file is not end
   2) We are reading these 6 variable in 3 aThing struct variable.
   3) Function theseAreGreenEggsAndHam checks if these three points are collinear, that is
       if slope of (A, C) is same as slope of (A, B)
       so, if (A.one - C.one) / (A.two - C.two) == (A.one - B.one) / (A.two - B.two)) return true;
            else return false.
       for the 1st input A(1, 2), B(2, 4), C(3, 6)
           (A.one - C.one) / (A.two - C.two) = (1 - 3) / (2 - 6) = 2 / 4 = 1 / 2
           (A.one - C.one) / (A.two - C.two) = (1 - 2) / (2 - 4) = 1 / 2
           equal so, it returns false,
       for the 2nd input A(-8, 2), B(0, 0), C(6, 3)
           (A.one - C.one) / (A.two - C.two) = (-8 - 6) / (2 - 3) = -14 / -1 = 14 / 1
           (A.one - C.one) / (A.two - C.two) = (-8 - 0) / (2 - 0) = -8/ 2 = -4 / 1
           not equal, so it returns false
   4) Print yes if lines are collinear, along with points
       Print no if lines are not collinear, along with points

   output:

   sanjiv@sanjiv-VirtualBox:~$ g++ file1.cpp -o file1
   sanjiv@sanjiv-VirtualBox:~$
   sanjiv@sanjiv-VirtualBox:~$ ./file1
   YES :1 2,2 4,3 6.
   NO :-8 2,0 0,6 3.